# Attosecond physics hidden in Cherenkov radiation

### v2

@@ Introduction
Close to the Cherenkov angle, the flash duration is directly tied to the electron coherence length upon entering the medium, opening avenues for contolling this duration by selecting packets of desired length and accelerating them in the low-current regime with no space-charge effects, typical for TEMs. Alternatively, measuring photon pulse durations can provide a novel method for determining electron packet length, a complementary approach to interferometry that is reminiscent of bunch length measurements via classical coherence [14]. The developed framework provides visualization of the emitted photon energy in real space and time, which is particularly desirable for biomedical applications of ChR in radiotherapy [23, 24] where only the classical Tamm-Frank theory in the far field has been used so far, apparently overestimating the process intensity in the formation zone. Naturally, ChR serves as an exemplary case within a broader class of media-induced emission phenomena where attosecond time scales are accessible via the phase-space analysis.
@@
Acknowledgment. We are grateful to D. Glazov, I. Ivanov, A. Shchepkin, V. Serbo, I. Pavlov, and especially to A. Tishchenko for useful discussions and suggestions, and also to Maria Zhuravleva for the help with the 3D picture. The studies on photon emission in phase space are supported by the Foundation for the Advancement of Theoretical Physics and Mathematics “BASIS”. Those on the temporal features of radiation at the finite distance by the Russian Science Foundation (Project No. 23-62-10026 https://rscf.ru/en/project/23-62-10026/). The analysis of the quantum shift is supported by the Government of the Russian Federation through the ITMO Fellowship and Professorship Program.

### v3

@@ Conclusions
The developed quantum framework provides visualization of the emitted photon energy in phase space and, in particular, in real space and time, including the near-field zone. This can be desirable for estimating the Cherenkov wakefields in accelerator chambers as well as for biomedical applications of ChR in radiotherapy [23, 24] where only the classical Tamm-Frank theory in the far field has been used so far, apparently overestimating the radiation intensity in the formation zone. We have predicted several spatiotemporal features of the quantum ChR, absent within the conventional momentum-space analysis limited to the far-field approximation. In particular, the Cherenkov flash duration close to the Cherenkov angle turns out to be determined by the electron coherence length upon entering the medium, opening avenues for controlling this duration by selecting packets of the desired length and accelerating them in the low-current regime with no space-charge effects, typical for TEMs. By using the charged particles with the Lorentz factors $\gamma=\varepsilon/m$ from $\gamma\sim 10$ to $\gamma\sim 10^{3}-10^{4}$, one can generate attosecond photon flashes and even the sub-attosecond ones if the target is installed close to the particle source. Such short photon pulses can come in handy for various studies in atomic physics, including those of the temporal effects in electron excitation and photoionization.
Alternatively, measuring photon pulse durations with subpicosecond resolution can provide a technique for determining the lengths of the emitting charged-particle packets. This can be used to control the lengths of the wave packets in electron microscopes and accelerator facilities (linacs or storage rings with electrons, protons, or ions), as well as of ultrarelativistic charged particles coming from the cosmos. Naturally, ChR serves as an exemplary case within a broader class of media-induced emission phenomena where attosecond time scales are accessible via phase-space analysis.
@@
Acknowledgment. We are grateful to D. Glazov, I. Ivanov, V. Serbo, I. Pavlov, and especially to A. Tishchenko for useful discussions and suggestions, and also to Maria Zhuravleva for the help with the 3D picture. The studies of general temporal features of Cherenkov radiation are supported by the Ministry of Science and Higher Education of the Russian Federation (Project FSER-2025-0012). Those of the photon field at a finite distance are supported by the Russian Science Foundation (Project No. 23-62-10026 https://rscf.ru/en/project/23-62-10026/). The analysis of the quantum shift is supported by the Government of the Russian Federation through the ITMO Fellowship and Professorship Program. The studies on emission of an electron packet in phase space (by D. Karlovets, D. Grosman, A. Shchepkin, and G. Sizykh) are supported by the Foundation for the Advancement of Theoretical Physics and Mathematics “BASIS”.
Author contributions. All authors contributed extensively to this work. Dm.K. conceived the idea, developed the theoretical framework, wrote the manuscript with input from all the authors and acquired the funding. D.G., G.S., A.C., A.Sh. and D.K. checked the calculations and interpreted the results in detail. D.G., A.C., and G.S. assisted with the figures, while A.C., Dm.K., and G.S. refined the presentation of results and theoretical details.
Data Availability. Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Competing interests. The authors declare no competing interests.